OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −1 −4 4 −3 3 @AccessDenied ?

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

3\2 *x =6/2 2/3*x=-1/3 so (3/2 +2/3)*x=3-1/3